# Netan Dogra

### v1

@@ 1. Introduction
For example, in [${}^{+}$08], Bugeaud, Mignotte, Siksek, Stoll and Tengely consider the genus two curve
@@ 1.1. $2$-descent for hyperelliptic curves
$H^{1}(\operatorname{Gal}(K^{\mathrm{sep}}|K),J[2])\simeq\mathrm{Ker}(K_{f}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}K^{\times}\otimes\mathbb{F}_{2}),$
@@
When $K$ is a number field, this isomorphism may be used to describe the Selmer group $\mathrm{Sel}_{2}(J)$ as a subgroup of $H^{1}(\operatorname{Gal}(K^{\mathrm{sep}}|K),J[2])$ (see [Sch95], [PS97], [Sto01]). This isomorphism is the basis of existing algorithms for proving upper bounds on the rank of the Jacobian of $X/K$, and is hence an essential tool in the implementation of the Chabauty–Coleman method for computing rational points on hyperelliptic curves whose Jacobians have small Mordell–Weil rank.
@@
The notation $H^{1}_{f}(\operatorname{Gal}(\mathbb{Q}_{S}|\mathbb{Q}),T\otimes\mathbb{F}_{p})$ is perhaps inadvisable, because in fact the subspace $H^{1}_{f}(G_{\mathbb{Q},S},T_{\otimes}\mathbb{F}_{p})$ will not be an invariant of $T\otimes\mathbb{F}_{p}$ but will depend on $T$. Recent work of Iovita and Marmora explores the extent to which the image of $H^{1}_{f}(G_{\mathbb{Q}_{p}},T)$ in $H^{1}(G_{\mathbb{Q}_{p}},T\otimes\mathbb{F}_{p})$ depends on the lift $T$ of $T\otimes\mathbb{F}_{p}$ [IM15]. In [Gaz18], Gazaki explores this in the case where $T$ is a tensor product of Tate modules of abelian varieties, which is essentially the case of interest for quadratic Chabauty. In these papers one can obtain a subspace depending only on $T\otimes\mathbb{F}_{p}$ if one imposes a condition on the Hodge–Tate weights, but since we are interested in $p=2$ in this paper we are working outside the range where their results apply.
One tool used for studying the Bloch–Kato Selmer group $H^{1}_{f}(\operatorname{Gal}(\overline{\mathbb{Q}}|\mathbb{Q}),\wedge^{2}V)$ is its relation to quotients of the étale fundamental group. More precisely, let $K|\mathbb{Q}$ be a field extension, and let $b$ be a point in $X(K)$. Let $\pi_{1}^{\mathrm{\acute{e}t},2}(X_{\overline{K}}-\infty,b)$ denote the maximal pro-$2$ quotient of the étale fundamental group of $X_{\overline{K}}-\infty$ with basepoint $b$. Let $\Pi_{2}$ denote its maximal $2$-nilpotent quotient. Then $\Pi_{2}$ is a central extension of $T_{2}(J)$ by $\wedge^{2}T_{2}(J)$ There is a unique quotient $\Gamma$ of $\Pi_{2}$ sitting in a commutative diagram with exact rows
@@
where the outer vertical maps are the obvious ones. The main results proved on “$\Gamma$-descent” for $X$ proved in this paper may be viewed as nonabelian generalisations of the main results described above for $2$-descent on hyperelliptic curves. Firstly, we give an explicit field-theoretic description of the cohomology group $H^{1}(\operatorname{Gal}(\overline{K}|K),\wedge^{2}J[2])$, in terms of units mod squares of a certain étale algebra of degree $2g^{2}+g$ over $K$. If $\operatorname{Gal}(\overline{K}|K)$ acts $2$-transitively on the roots of the polynomial $f$, then this étale algebra is a field extension of $K$.
Given a pair of points $z,b$ in $X(K)$, we obtain a class $\kappa_{\Gamma,b}(z)$ in $H^{1}(\operatorname{Gal}(\overline{K}|K),\wedge^{2}J[2])$ whose image in $H^{1}(\operatorname{Gal}(\overline{K}|K),J[4])$ is the mod $4$ Kummer class $\kappa_{4,b}(z)$ of $[z]-[b]$. If $[z]-[b]$ lies in $4\cdot\mathrm{Jac}(X)(K)$, there is an element of $H^{1}(\operatorname{Gal}(\overline{K}|K),\wedge^{2}J[2])$ with image $\kappa_{\Gamma,b}(z)$ in $H^{1}(\operatorname{Gal}(\overline{K}|K),\Gamma)$. We give an explicit formula for such a class in the case when $X$ is hyperelliptic with a rational Weierstrass point, and the points $z$ and $b$ are not Weierstrass points.
@@
$K[x,y,\frac{1}{x-y}]/(f(x),f(y)),$
and define $K_{f}^{(2)}:=K_{f,2}^{\tau}$, where $\tau$ is the involution sending $x$ to $y$. Let $\alpha$ and $\beta$ denote the image of $x$ and $y$ in $K_{f,2}$.
@@
    $H^{1}(\operatorname{Gal}(\overline{K}|K),\wedge^{2}J[2])\simeq\mathrm{Ker}(L_{\mathrm{sym}}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}K(\alpha)^{\times}\otimes\mathbb{F}_{2}).$
@@
    $\left(1+\frac{c_{\alpha}u_{\alpha}-c_{\beta}u_{\beta}}{\alpha-\beta}\right)$
@@
In the sequel to this paper [Dog23], we study liftability or ‘second descent’ obstructions for $H^{1}_{f}(G_{\mathbb{Q}_{v}},\wedge^{2}T_{2}J)$, and use these to verify finiteness of $X(\mathbb{Q}_{2})_{2}$ for several thousand genus 2 curves with Jacobians of Mordell–Weil rank 2 and Picard number 1.
@@ 1.2. Notation
For a number field $K$ and a finite set $S$ of places of $K$, we define $\mathrm{Cl}(\mathcal{O}_{K,S}):=\mathrm{Pic}(\mathcal{O}_{K,S})$. For any integer $N>0$, we define $(L^{\times}\otimes\mathbb{Z}/N\mathbb{Z})_{S}$ to be the subgroup of $L^{\times}\otimes\mathbb{Z}/N\mathbb{Z}$ consisting of elements whose image in $L_{v}^{\times}\otimes\mathbb{Z}/N\mathbb{Z}$ lies in the image of $\mathcal{O}_{v}\otimes\mathbb{Z}/N\mathbb{Z}$ for all $v$ not in $S$.
@@ Acknowledgements
I am grateful to Ed Schaefer for explaining $2$-Selmer groups of hyperelliptic curves to me, and to Michael Stoll for several detailed suggestions about an earlier version of this work. This research was supported by a Royal Society University Research Fellowship.
@@ 2. Galois cohomology of the wedge square
In this section, unless otherwise indicated, $K$ will denote a field of characteristic different from $2$ and $f\in K[x]$ will be a monic separable polynomial of degree $d$.
@@
Let $G$ be a group, and let $S$ be a finite $G$-set, and let $k$ be a field of characteristic $2$. Then we have an isomorphism of $G$-modules
@@
$K_{f}^{(2),\times}\otimes\mathbb{F}_{2}\to K_{f,2}^{\times}\otimes\mathbb{F}_{2}\oplus K_{f}\otimes\mathbb{F}_{2}$
@@
(5) $0\to J[2]\to\mathrm{Ind}^{K_{f}}_{K}\mathbb{F}_{2}\to\mathbb{F}_{2}$
@@
On the other hand the splitting of (5) tells us that this isomorphic to $\mathrm{Ker}(H^{1}(K,\wedge^{2}\mathrm{Ind}^{K_{f}}_{K}\mathbb{F}_{2})\to H^{1}(K,\mathrm{Ind}^{K_{f}}_{K}\mathbb{F}_{2}))$. Here the map is equal to the map on cohomology induced by
@@ 3. The nonabelian $(x-T)$-map
$(X-\{\infty\})(K)\stackrel{{\scriptstyle j_{X-\infty,b}}}{{\longrightarrow}}H^{1}(K,\pi_{1}^{\mathrm{\acute{e}t}}(X_{K^{\mathrm{sep}}}-\{\infty\},b))$
which may be defined by sending a point $z\in X-\infty(K)$ to the Galois equivariant continuous $\pi_{1}^{\mathrm{\acute{e}t}}(X_{K^{\mathrm{sep}}}-\infty,b)$-torsor of paths $\pi_{1}^{\mathrm{\acute{e}t}}(X_{K^{\mathrm{sep}}}-\infty;b,z)$ (and similarly for $X-\{\infty\}$). More generally, for any geometrically connected scheme $Z$ over $K$, and $b\in Z(K)$, and any $\operatorname{Gal}(K^{\mathrm{sep}}|K)$-stable quotient $G$ of $\pi_{1}^{\mathrm{\acute{e}t}}(Z_{K^{\mathrm{sep}}},b)$, we define a map
@@
For example, let $\Gamma$ denote the quotient from the introduction(defined in (1)), and let $\kappa_{\Gamma,b}(z)$ in $H^{1}(\operatorname{Gal}(\overline{K}|K),\Gamma)$ denote the section class associated to $b$ and $z$. Then the image of $\kappa_{\Gamma,b}(z)$ in $H^{1}(\operatorname{Gal}(\overline{K}|K),J[4])$ is the mod $4$ Kummer class $\kappa_{4,b}(z)$ of $[z]-[b]$. If $[z]-[b]$ lies in $4\cdot\mathrm{Jac}(X)(K)$, there is an element of $H^{1}(\operatorname{Gal}(\overline{K}|K),\wedge^{2}J[2])$ with image $\kappa_{\Gamma,b}(z)$ in $H^{1}(\operatorname{Gal}(\overline{K}|K),\Gamma)$.
@@
    If the kernel of $\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{K}},b)\to G$ is stable under the conjugation action of $\operatorname{Gal}(\overline{K}|K)$, then there is a descent $(Y^{\prime},c)$ of $Y$ to a pointed cover $(Y^{\prime},c)\to(X,b)$ defined over $K$.
@@
It is enough to prove this after base change to $K^{\mathrm{sep}}$. The function field of the covering corresponding to $J[2]$ is generated over $K^{\mathrm{sep}}(X)$ by $\{u_{\alpha}:\alpha\in\mathrm{Roots}(f)\}$. We have a norm map
@@
Base changing to $K^{\mathrm{sep}}$ and applying Lemma 9 we see that $X_{2}$ decomposes into curves $X_{2}^{+}$ and $X_{2}^{-}$ defined by $\mathrm{Nm}(u_{\alpha})=y$ and $\mathrm{Nm}(u_{\alpha})=-y$ respectively. ∎
@@
Let $K^{\mathrm{sep}}(u_{\alpha},u_{\beta},w_{\alpha\beta})$ be the field extension generated by $w_{\alpha\beta}$ satisfying
$w_{\alpha\beta}^{2}:=2\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\beta-\alpha}\right).$
@@
$K^{\mathrm{sep}}(X)(u_{\alpha},u_{\beta},z_{\alpha\beta})=K^{\mathrm{sep}}(u_{\alpha},u_{\beta},w_{\alpha\beta}),$
where
(10) $w_{\alpha\beta}^{2}:=2\left(1+\frac{c_{\alpha}u_{\alpha}-c_{\beta}u_{\beta}}{\beta-\alpha}\right).$
@@
By Lemma 8, the class of $\frac{u_{\alpha}-\sqrt{(\beta-\alpha)/c_{\alpha}}}{1-\sqrt{(\beta-\alpha)/c_{\alpha}}}$ in $K(\alpha,\beta,\sqrt{(\beta-\alpha)/c_{\alpha}})(Y_{\alpha}\times_{X}Y_{\beta})$ is equal to that of
$2(c_{\alpha}+(\beta-\alpha))/c_{\beta}-u_{\beta},$
since
$\mathrm{Tr}\left(\frac{u_{\alpha}-\sqrt{(\beta-\alpha)/c_{\alpha}}}{1-\sqrt{(\beta-\alpha)/c_{\alpha}}}\right)$ $=2\frac{c_{\alpha}u_{\alpha}+\beta-\alpha}{c_{\beta}}$
@@
where the trace and norm are from $K(\alpha,\beta,\sqrt{(\beta-\alpha)/c_{\alpha}})(Y_{\alpha}\times_{X}Y_{\beta})$ to $K(\alpha,\beta)(Y_{\alpha}\times_{X}Y_{\beta})$. On the other hand the class of $2(c_{\alpha}+(\beta-\alpha))/c_{\beta}-u_{\beta}$ is equal to that of (10). ∎

### v2

@@ 1. Introduction
For example, in [BMS+08], Bugeaud, Mignotte, Siksek, Stoll and Tengely consider the genus two curve
@@ 1.1. $2$-descent for hyperelliptic curves
$H^{1}(K,J[2])\simeq\mathrm{Ker}(K_{f}^{\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}K^{\times}\otimes\mathbb{F}_{2}),$
@@
When $K$ is a number field, this isomorphism may be used to describe the Selmer group $\mathrm{Sel}_{2}(J)$ as a subgroup of $H^{1}(K,J[2])$ (see [Sch95], [PS97], [Sto01]). This isomorphism is the basis of existing algorithms for proving upper bounds on the rank of the Jacobian of $X/K$, and is hence an essential tool in the implementation of the Chabauty–Coleman method for computing rational points on hyperelliptic curves whose Jacobians have small Mordell–Weil rank.
@@
The notation $H^{1}_{f}(\operatorname{Gal}(\mathbb{Q}_{S}|\mathbb{Q}),T\otimes\mathbb{F}_{p})$ is perhaps inadvisable, because in fact the subspace $H^{1}_{f}(G_{\mathbb{Q},S},T\otimes\mathbb{F}_{p})$ will not be an invariant of $T\otimes\mathbb{F}_{p}$ but will depend on $T$. Recent work of Iovita and Marmora explores the extent to which the image of $H^{1}_{f}(\mathbb{Q}_{p},T)$ in $H^{1}(\mathbb{Q}_{p},T\otimes\mathbb{F}_{p})$ depends on the lift $T$ of $T\otimes\mathbb{F}_{p}$ [IM15]. In [Gaz18], Gazaki explores this in the case where $T$ is a tensor product of Tate modules of abelian varieties, which is essentially the case of interest for quadratic Chabauty. In these papers one can obtain a subspace depending only on $T\otimes\mathbb{F}_{p}$ if one imposes a condition on the Hodge–Tate weights, but since we are interested in $p=2$ in this paper we are working outside the range where their results apply.
One tool used for studying the Bloch–Kato Selmer group $H^{1}_{f}(\operatorname{Gal}(\overline{\mathbb{Q}}|\mathbb{Q}),\wedge^{2}V)$ is its relation to quotients of the étale fundamental group. More precisely, let $K|\mathbb{Q}$ be a field extension, and let $b$ be a point in $X(K)$. Let $\pi_{1}^{\mathrm{\acute{e}t},2}(X_{\overline{K}}-\{\infty\},b)$ denote the maximal pro-$2$ quotient of the étale fundamental group of $X_{\overline{K}}-\{\infty\}$ with basepoint $b$. Let $\Pi_{2}$ denote its maximal $2$-nilpotent quotient. Then $\Pi_{2}$ is a central extension of $T_{2}(J)$ by $\wedge^{2}T_{2}(J)$. There is a unique quotient $\Gamma$ of $\Pi_{2}$ sitting in a commutative diagram with exact rows
@@
where the outer vertical maps are the obvious ones. The main results on “$\Gamma$-descent” for $X$ proved in this paper may be viewed as nonabelian generalisations of the main results described above for $2$-descent on hyperelliptic curves. Firstly, we give an explicit field-theoretic description of the cohomology group $H^{1}(K,\wedge^{2}J[2])$, in terms of units mod squares of a certain étale algebra of degree $2g^{2}+g$ over $K$. If $\operatorname{Gal}(\overline{K}|K)$ acts $2$-transitively on the roots of the polynomial $f$, then this étale algebra is a field extension of $K$.
Given a pair of points $z,b$ in $X(K)$, we obtain a class $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,\Gamma)$ whose image in $H^{1}(K,J[4])$ is the mod $4$ Kummer class $\kappa_{4,b}(z)$ of $[z]-[b]$. If $[z]-[b]$ lies in $4\cdot\mathrm{Jac}(X)(K)$, there is an element of $H^{1}(K,\wedge^{2}J[2])$ with image $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,\Gamma)$. We give an explicit formula for such a class in the case when $X$ is hyperelliptic with a rational Weierstrass point, and the points $z$ and $b$ are not Weierstrass points.
@@
$K[s,t,\frac{1}{s-t}]/(f(s),f(t)),$
and define $K_{f}^{(2)}:=K_{f,2}^{\tau}$, where $\tau$ is the involution sending $s$ to $t$. Let $\alpha$ and $\beta$ denote the image of $s$ and $t$ in $K_{f,2}$.
@@
    $H^{1}(K,\wedge^{2}J[2])\simeq\mathrm{Ker}(K_{f}^{(2),\times}\otimes\mathbb{F}_{2}\stackrel{{\scriptstyle\mathrm{Nm}}}{{\longrightarrow}}K_{f}^{\times}\otimes\mathbb{F}_{2}).$
@@
    $2\left(1+\frac{c_{\alpha}u_{\alpha}-c_{\beta}u_{\beta}}{\beta-\alpha}\right)$
@@
In the sequel to this paper [Dog23], we study liftability or ‘second descent’ obstructions for $H^{1}_{f}(\mathbb{Q}_{v},\wedge^{2}T_{2}J)$, and use these to verify finiteness of $X(\mathbb{Q}_{2})_{2}$ for several thousand genus 2 curves with Jacobians of Mordell–Weil rank 2 and Picard number 1.
@@ 1.2. Notation
For a number field $K$ and a finite set $S$ of places of $K$, we define $\mathrm{Cl}(\mathcal{O}_{K,S}):=\mathrm{Pic}(\mathcal{O}_{K,S})$. For any integer $N>0$, we define $(K^{\times}\otimes\mathbb{Z}/N\mathbb{Z})_{S}$ to be the subgroup of $K^{\times}\otimes\mathbb{Z}/N\mathbb{Z}$ consisting of elements whose image in $K_{v}^{\times}\otimes\mathbb{Z}/N\mathbb{Z}$ lies in the image of $\mathcal{O}_{v}\otimes\mathbb{Z}/N\mathbb{Z}$ for all $v$ not in $S$.
@@ Acknowledgements
I am grateful to Ed Schaefer for explaining $2$-Selmer groups of hyperelliptic curves to me, and to Michael Stoll for several detailed suggestions about an earlier version of this work, and corrections and comments on an earlier version of this paper. I am grateful to Lee Berry for many corrections to a previous version of this paper and useful discussions. This research was supported by a Royal Society University Research Fellowship.
@@ 2. Galois cohomology of the wedge square
In this section, unless otherwise indicated, $K$ will denote a field of characteristic different from $2$ and $f\in K[x]$ will be a monic separable polynomial of degree $d$. None of our general results on étale algebras and Galois cohomology are original (see for example [BPS16] or [SS04]), but we include them for the sake of completeness and to establish notation.
@@
Let $G$ be a group, let $S$ be a finite $G$-set, and let $k$ be a field of characteristic $2$. Then we have an isomorphism of $G$-modules
@@
$K_{f}^{(2),\times}\otimes\mathbb{F}_{2}\to K_{f,2}^{\times}\otimes\mathbb{F}_{2}\times K_{f}\otimes\mathbb{F}_{2}$
@@
(5) $0\to J[2]\to\mathrm{Ind}^{K_{f}}_{K}\mathbb{F}_{2}\to\mathbb{F}_{2}\to 0$
@@
On the other hand the splitting of (5) tells us that this is isomorphic to $\mathrm{Ker}(H^{1}(K,\wedge^{2}\mathrm{Ind}^{K_{f}}_{K}\mathbb{F}_{2})\to H^{1}(K,\mathrm{Ind}^{K_{f}}_{K}\mathbb{F}_{2}))$. Here the map is equal to the map on cohomology induced by
@@ 3. The nonabelian $(x-T)$-map
$(X-\{\infty\})(K)\stackrel{{\scriptstyle j_{X-\{\infty\},b}}}{{\longrightarrow}}H^{1}(K,\pi_{1}^{\mathrm{\acute{e}t}}(X_{K^{\mathrm{sep}}}-\{\infty\},b))$
which may be defined by sending a point $z\in X-\{\infty\}(K)$ to the Galois equivariant continuous $\pi_{1}^{\mathrm{\acute{e}t}}(X_{K^{\mathrm{sep}}}-\{\infty\},b)$-torsor of paths $\pi_{1}^{\mathrm{\acute{e}t}}(X_{K^{\mathrm{sep}}}-\{\infty\};b,z)$ (and similarly for $X-\{\infty\}$). More generally, for any geometrically connected scheme $Z$ over $K$, and $b\in Z(K)$, and any $\operatorname{Gal}(K^{\mathrm{sep}}|K)$-stable quotient $G$ of $\pi_{1}^{\mathrm{\acute{e}t}}(Z_{K^{\mathrm{sep}}},b)$, we define a map
@@
For example, let $\Gamma$ denote the quotient defined in (1), and let $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,\Gamma)$ denote the section class associated to $b$ and $z$. Then the image of $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,J[4])$ is the mod $4$ Kummer class $\kappa_{4,b}(z)$ of $[z]-[b]$. If $[z]-[b]$ lies in $4\cdot\mathrm{Jac}(X)(K)$, there is an element of $H^{1}(K,\wedge^{2}J[2])$ with image $\kappa_{\Gamma,b}(z)$ in $H^{1}(K,\Gamma)$.
@@
Suppose the kernel of $\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{K}},b)\to G$ is stable under the conjugation action of $\operatorname{Gal}(\overline{K}|K)$.
@@
    There is a descent $(Y^{\prime},c)$ of $Y$ to a pointed cover $(Y^{\prime},c)\to(X,b)$ defined over $K$.
@@
We have a norm map
@@
We claim that $X_{2}$ is a disjoint union of connected components $X_{2}^{+}=\{\mathrm{Nm}(u_{\alpha})\}=y$ and $X_{2}^{-}=\{\mathrm{Nm}(u_{\alpha})=-y\}$. It is enough to prove this after base change to $K^{\mathrm{sep}}$. The function field of the covering corresponding to $J[2]$ is generated over $K^{\mathrm{sep}}(X)$ by $\{u_{\alpha}:\alpha\in\mathrm{Roots}(f)\}$. The result then follows from base changing to $K^{\mathrm{sep}}$ and applying Lemma 9. ∎
@@
Let $K^{\mathrm{sep}}(X)(u_{\alpha},u_{\beta},w_{\alpha\beta})$ be the field extension of $K^{\mathrm{sep}}(X)(u_{\alpha},u_{\beta})$ generated by $w_{\alpha\beta}$ satisfying
(10) $w_{\alpha\beta}^{2}:=2\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\beta-\alpha}\right).$
@@
$K^{\mathrm{sep}}(X)(u_{\alpha},u_{\beta},z_{\alpha\beta})=K^{\mathrm{sep}}(X)(u_{\alpha},u_{\beta},w_{\alpha\beta}).$
@@
We have
$\mathrm{Tr}\left(\frac{u_{\alpha}-\sqrt{(\beta-\alpha)/c_{\alpha}}}{1-\sqrt{(\beta-\alpha)/c_{\alpha}}}\right)$ $=2\frac{c_{\alpha}u_{\alpha}+\alpha-\beta}{c_{\beta}}$
@@
where the trace and norm are from $K(\alpha,\beta,\sqrt{(\beta-\alpha)/c_{\alpha}})(Y_{\alpha}\times_{X}Y_{\beta})$ to $K(\alpha,\beta)(Y_{\alpha}\times_{X}Y_{\beta})$. Hence, by Lemma 8, the class of $\frac{u_{\alpha}-\sqrt{(\beta-\alpha)/c_{\alpha}}}{1-\sqrt{(\beta-\alpha)/c_{\alpha}}}$ in $K(\alpha,\beta,\sqrt{(\beta-\alpha)/c_{\alpha}})(Y_{\alpha}\times_{X}Y_{\beta})^{\times}\otimes\mathbb{F}_{2}$ is equal to that of
$\frac{2(\alpha-\beta)}{c_{\beta}}\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\alpha-\beta}\right).$
This is equal to $2\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\alpha-\beta}\right)$ in $K^{\mathrm{sep}}(Y_{\alpha}\times_{X}Y_{\beta})^{\times}\otimes\mathbb{F}_{2}$. Finally, note that
$\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\alpha-\beta}\right)\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\beta-\alpha}\right)$
$=$ $-c_{\alpha}c_{\beta}\left(\frac{(u_{\beta}-1)(u_{\alpha}+1)}{(\alpha-\beta-c_{\beta}u_{\beta}-c_{\alpha}u_{\alpha})}\right).$
Hence
$2\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\alpha-\beta}\right)=\left(1+\frac{c_{\alpha}u_{\alpha}(z)-c_{\beta}u_{\beta}(z)}{\beta-\alpha}\right)$
in $K^{\mathrm{sep}}(Y_{\alpha}\times_{X}Y_{\beta})^{\times}\otimes\mathbb{F}_{2}$. ∎